# Count of Monte Crypto: Accounting-based Defenses for Cross-Chain Bridges

### v1

@@ 8 Discussion and Future Work
This paper is a “first cut” at identifying, validating and implementing a generic mode of theft protection into cross-chain transactions. We show that a simple balance invariant is sufficient to detect and prevent most bridge attacks. Our effort also highlights the need for better transparency and auditing in the bridge ecosystem, and for new bridge designs. Indeed, implementing and deploying a cross-chain token bridge with an in-line balance invariant monitor is an open challenge.
@@
As mentioned in Section 2, in this paper we focus on cross-chain token bridges and thus our approach does not directly apply to bridges that have built-in exchanges and swaps. Still, similar invariant-checking techniques likely extend to such cross-chain exchanges and similar protocols (e.g., Automated Market Makers which trade between a range of cryptocurrencies in decentralized exchanges). Such extensions will require incorporating contemporaneous price oracle data into our framework and reasoning about the variation therein.
@@
The larger opportunities opened up by this work are more architectural. Today, when a crypto platform declares that their system has been audited, they typically are referring to a third-party that inspects their smart contracts for common logic errors or deficiencies in their process. Any validation of overall financial safety is inherently entwined with the totality of the implementation. However, such code-oriented audits are inherently challenging and ill-suited for overarching questions of financial risk. They must evaluate all contracts and bridge code that directly or indirectly have access to keys controlling assets and then must ensure that no combination of inputs or invocations might lead to a loss (itself, an outcome that is rarely well defined a priori). Unsurprisingly, multiple bridges in our study had been previously audited by third parties before their attacks.
@@
Our work suggests that by separating overall financial safety invariants from the intricate details of every given contract or trade, we can dramatically simplify the defender’s job and their ability to reason about risk. Similar to so-called “circuit breakers” implemented in traditional securities exchanges, or the atomic repayment guarantees embedded in single-blockchain “flash loan” services, the ability to script overarching financial constraints independent of transaction details can be extremely powerful. For example, one might extend this work beyond a simple cross-chain balance invariant to enforce limits on liquidity losses or to place limits on collateralization risk for tokens backed by stablecoins. We hypothesize that by centralizing the specification and enforcement of overall financial constraints in one place, they will be far easier to audit and reason about.
Open Science Policy
We will make all of our data, results, and code public upon publication. In addition, we have already included some of the most important results in the appendix.
Ethics
@@
We believe our work, which deals with public data, no identified individuals and a simple means for identifying attacks on crypto token transfer bridges (and potentially preventing such attacks in the future), has very low ethical risk and significant upside. However, in the interest of due diligence, we provide a more thorough justification of this position following the Menlo Report principles as indicated by the USENIX Security call.
Considering the Menlo report principle of “Respect for Persons”, we are not dealing with, nor do we attempt to analyze, data at the granularity of individual persons. We strictly use public data and to the extent that we identify outcomes (e.g., attacks) they are at the granularity of bridges and blockchains, not individuals. Similarly, with respect to the principle of “Justice”, we have no reason to believe that our analysis discriminates at the level of persons or distinct groups of persons — all users of bridges should benefit equally (excepting criminals). On the topic of “Respect for Law and Public Interest”, we are engaging with existing public data, we have documented our methods clearly and we are careful to limit our factual statements to those for which we have evidence. As well, we have attempted to disclose significant suspicious bridge transactions to appropriate bridge operators (yet with limited success as most of the bridges have ceased operations).
Finally, on the topic of “Beneficence”, we identify the three sets of stakeholders in our work. These include the past and (potential) future victims whose investments, both direct and indirect, via token transfer bridges have been undermined by large-scale thefts from these services. In addition, other stakeholders include bridge operators and, finally, the criminal actors who derive income from vulnerabilities in bridges. We argue that our retrospective analysis creates no direct new harm for any parties — these attacks have already happened. For bridge operators, our analysis may create some potential liability in supporting a civil claim of poor stewardship, but at the same time our documentation of the widespread nature of this problem provides a “common practice” defense. Some of the unilateral transfers we have documented, if they were illegitimate (which we do not know), could also create liability for bridge operators, but we believe this potential harm is balanced against the interest of investors in transparency and fairness in operating financial services. Finally, for our proposed real-time analysis approaches, if these were adopted by bridge operators, it could foreclose a broad range of attacks. We admit that this outcome could harm our criminal stakeholders as it might deny them of an important income stream. Knock on effects could include a reduction in revenues for online crypto money laundering services and Lamborghini dealerships, but we believe this is well-justified by the concordant reduction in victimization.
Appendix A Attack Descriptions and Additional Details
@@ A.1 Ronin Bridge
A.2 Poly Network Bridge (2021)
@@ A.3 Binance Token Hub
Results. CrossChecked analyzed over 2M bridge transactions between Binance Beacon Chain and Binance Smart Chain. CrossChecked only alerted on the two attack transactions, citing the discrepancy between the amount given out and the amount received by the bridge. We end by noting that the partner deposit transactions returned by the Binance Token Hub’s API for two attack transactions are different than the ones suggested in some blogposts [8]. However, the attack transactions will be flagged regardless of which partner deposit transactions is used.
A.4 Wormhole
@@ A.5 Nomad Bridge
Results. CrossChecked analyzed over 37k transactions, and alerted on 962 transactions that transferred assets to 561 unique addresses. We found one dataset that reported 561 addresses, which matched exactly with the addresses we identified. We also found one Github repository that mentioned identifying 960 transactions [20]. As mentioned earlier, we identified two more transactions. Upon manually inspection, we confirm that the two transactions were indeed malicious.
A.6 Harmony Bridge
@@ A.9 Anyswap Bridge
A.10 Poly Network Bridge (2023)
@@ A.12 Meter Bridge
Wormhole #1 0x542efe5a7f965b927a294bce7f3a30492d0e2ca60da2e2d6cdfaf523a28a2a9c Polygon 0.5 WSOL 5AiesW9pKrZvCCJM8QPWmqx...dkadV1waPVLWfsnCMVQmyYaciLxEo SOL 650 USDC
Wormhole #2 0x8bbb7befd198a5e90297f451fc43a9e90de083289a041c8af94116c785cf496d Polygon 4.3k MATIC 0xbe81895b1c3172fd69b8d4d9bf726edfdc17083876c440f9414ff316999237d7 AVAX 2.3k WMATIC
@@
Table 3: Other flagged transactions. PolyNet stands for Poly Network. Full transaction hash for Solana: 5AiesW9pKrZvCCJM8QPWmqxsnRoQwHaQmX8NR9a8BFz3pmt2ypW67zgqeRWdkadV1waPVLWfsnCMVQmyYaciLxEo.
Appendix B Additional Transactions Identified

### v3

@@ 9. Discussion and Future Work
This paper is a “first cut” at identifying, validating and implementing a generic mode of theft protection into cross-chain transactions. We show that a simple balance invariant is sufficient to detect and prevent most bridge attacks. Our effort also highlights the need for better auditability and auditing in the bridge ecosystem, and for new bridge designs.
@@
As mentioned in Section 2, in this paper we focus on cross-chain token bridges and thus our approach does not directly apply to bridges that have built-in exchanges and swaps. Still, similar invariant-checking techniques likely extend to such cross-chain exchanges and similar protocols (e.g., Automated Market Makers which trade between cryptocurrencies). Such extensions will require incorporating additional logic.
@@
The larger opportunities opened up by this work are more architectural. Today, when a crypto platform declares that their system has been audited, they typically are referring to a third-party that inspects their smart contracts for common logic errors or deficiencies in their process. Any validation of overall financial safety is inherently entwined with the totality of the implementation. However, such code-oriented audits are inherently challenging and ill-suited for overarching questions of financial risk. They must evaluate all contracts and code continuously and also ensure that no combination of inputs or invocations might lead to a loss. Unsurprisingly, multiple bridges in our study had been previously audited by third parties before their attacks.
@@
Our work suggests that by separating overall financial safety invariants from the intricate details of every given contract or trade, we can dramatically simplify the defender’s job and their ability to reason about risk. Similar to so-called “circuit breakers” implemented in traditional securities exchanges, the ability to script overarching financial constraints independent of transaction details can be extremely powerful. For example, one might extend this work beyond a simple cross-chain balance invariant to enforce limits on liquidity losses. We hypothesize that by centralizing the specification and enforcement of overall financial constraints in one place, they will be far easier to audit and reason about.
Appendix A Ethics
@@
We believe our work, which deals with public data, no identified individuals and a simple means for identifying attacks on crypto token transfer bridges (and potentially preventing such attacks in the future), has very low ethical risk and significant upside. Moreover, we have attempted to disclose significant suspicious bridge transactions to appropriate bridge operators (yet with limited success as most of the bridges have ceased operations). Lastly, we will make our code and data available upon publication.
Appendix B Attack Descriptions and Additional Details
@@ B.1. Ronin Bridge
B.2. Poly Network Bridge (2021)
@@ B.3. Binance Token Hub
Results. CrossChecked analyzed over 2M bridge transactions between Binance Beacon Chain and Binance Smart Chain. CrossChecked only alerted on the two attack transactions, citing the discrepancy between the amount given out and the amount received by the bridge. We end by noting that the partner deposit transactions returned by the Binance Token Hub’s API for two attack transactions are different than the ones suggested in some blogposts (emiliano, 2024). However, the attack transactions will be flagged regardless of which partner deposit transactions is used.
B.4. Wormhole
@@ B.5. Nomad Bridge
Results. CrossChecked analyzed over 37k transactions, and alerted on 962 transactions that transferred assets to 561 unique addresses. We found one dataset that reported 561 addresses, which matched exactly with the addresses we identified. We also found one Github repository that mentioned identifying 960 transactions (Nomad, ). As mentioned earlier, we identified two more transactions. Upon manually inspection, we confirm that the two transactions were indeed malicious.
B.6. Harmony Bridge
@@ B.9. Anyswap Bridge
B.10. Poly Network Bridge (2023)
@@ Appendix C Additional Transactions Identified
Wormhole #1 0x8bbb7befd198a5e90297f451fc43a9e90de083289a041c8af94116c785cf496d Polygon 0.5 WSOL 5AiesW9pKrZvCCJM8QPWmqx...dkadV1waPVLWfsnCMVQmyYaciLxEo SOL 650 USDC
Wormhole #2 0x55d1e486a8e2102e07fd6270a03f05bbee7b43bf27ebac97b95b98e068f6740e Polygon 4.3k MATIC 0xbe81895b1c3172fd69b8d4d9bf726edfdc17083876c440f9414ff316999237d7 AVAX 2.3k WMATIC
@@
Table 3. Other flagged transactions. PolyNet stands for Poly Network 2023. Full transaction hash for Solana: 5AiesW9pKrZvCCJM8QPWmqxsnRoQwHaQmX8NR9a8BFz3pmt2ypW67zgqeRWdkadV1waPVLWfsnCMVQmyYaciLxEo.